# Wenfan Ou and Sheng Bi

### v3

@@ 3 Parametric RMDPs
where $\Theta$ is a subset of $\mathbb{R}^{q}$ and $P^{\theta}(\cdot\mid s,a),~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ is an affine function from $\Theta$ to $[0,1]$ that satisfies $P^{\theta}(\cdot\mid s,a):=k_{sa}+K_{sa}\theta$ for some $k_{sa}\in\mathbb{R}^{|\mathcal{S}|}$ and $K_{sa}\in\mathbb{R}^{|\mathcal{S}|\times k}$. Note that $\theta$ controls the ambiguity sets essentially, which allows to condense all ambiguous information and parameters in the set $\Theta$, e.g.,
@@
where $O_{l}$ is a $k\times k$ matrix, $O_{l}\preceq 0$, $\omega$ is the constant and $L$ is the number of constraints applied to $\theta$. The benefit of adhering to this construction as defined in (8) is that the set $\Theta$ can be derived as the finite intersection of closed half-spaces and ellipsoids, enabling efficient policy iteration by tailoring to second-order cone programs and semi-definite programs. They assume that $\Theta$ is bounded and has a nonempty interior, which means none of the parameters in $\Theta$ is fully explained by the others. With the $s$-rectangular assumption, the authors show that solving an RMDP is equivalent to finding a fixed point of the robust Bellman operator, and provide the time complexity of policy evaluation and policy improvement, respectively.
@@
Black et al. (2022) investigate RMDPs with the ambiguity sets which only contain distribution from the same parametric family, focusing on identifying the worst-case parameters instead of the entire distribution. This approach aims to alleviate challenges with moment estimation in non-parametric RMDPs, particularly the issue of overly conservative solutions that can arise when parametric distributions are fitted well enough. For each $(s,a)\in\mathcal{S}\times\mathcal{A}$, $\mathbf{p}_{sa}$ is uniquely defined by the distribution of some exogenous random variable $\xi_{sa}\in\Xi_{sa}$. Let $f_{\xi_{sa}}$ be the probability mass function of $\xi_{sa}$ which is controlled by the parameters $\bm{\theta}_{sa}=(\theta_{sa1},\cdots,\theta_{sak})$. The authors assume that the next state $s_{t+1}$ is specified by a simple, known function $g$ of the exogenous random variable $\xi_{sa}$ as $s_{t+1}=g(\xi_{sa}|s,a)$. Hence, the transition probability $p_{sas^{\prime}}$ for each transition tuple $(s,a,s^{\prime})$ can be computed as
$p_{sas^{\prime}}=P^{\bm{\theta}_{sa}}(s^{\prime}|s,a)=P^{\bm{\theta}_{sa}}(g(\xi_{sa}|s,a)=s^{\prime})=\sum_{\xi\in\Xi_{sa}(s^{\prime})}f_{\xi_{sa}}(\xi|\bm{\theta}_{sa}),$
@@
where $\Xi_{sa}(s^{\prime})$ is the support of $\xi$ with $(s,a,s^{\prime})$ transition. Because the transition probability $\mathbf{p}_{sa}$ is uniquely specified by $\bm{\theta}$, the authors simplify ambiguity sets for $\bm{\theta}$, rather than $\mathbf{p}_{sa}$, to maintain the structure of $\mathbf{p}_{sa}$. More specifically, they consider ambiguity sets with $s$-rectangular structure $\Theta=\bigotimes_{s\in\mathcal{S}}\Theta_{s}$ where $\Theta_{s}\subseteq\mathbb{R}^{k},\mathrm{~{}}\forall s\in\mathcal{S}$, and then reformulate the RMDP as:
$\max_{\pi}\min_{\bm{\theta}\in\Theta}\mathbb{E}_{\mathbf{p}sa}^{\pi}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\right|\bm{\theta}\right].$ (9)
For $s\in\mathcal{S}$, $\Theta_{s}$ results from confidence regions, assuming $N$ samples from the true measure $P^{0}$ are accessible. By creating an MLE $\bm{\hat{\theta}}$ of the underlying $\bm{\theta}^{0}$, an approximate $1-\alpha$ confidence set for $\bm{\theta}_{s}$ is given by:
@@
$\Theta_{s}=\left\{\bm{\theta}_{s}\in\mathbb{R}^{|\mathcal{A}|}\times\mathbb{R}^{k}:\sum_{a\in\mathcal{A}}\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)^{T}I_{\mathbb{E}}\left(\hat{\bm{\theta}}_{sa}\right)\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)\leq\chi_{k\mathcal{A},1-\alpha}^{2}\right\}.$ (10)
Taking the non-linearity of the $P$ as functions of the parameters into consideration, the authors use a discretization $\Theta_{s}^{\prime}$ of the ambiguity set $\Theta_{s}$, and reformulate the Bellman update as a linear program with $|\Theta_{s}^{\prime}|+1$ constraints.
@@ 4 Moment-Based RMDPs
While the ideas of parametric RMPDs are natural, it is challenging in practice to determine rational parameter values range. Moreover, parametric approaches impose constraints on each probability that may be too restrictive, often resulting in overly conservative or unrealistic solutions. On the other hand, it is reasonable to assume that the decision-maker can access historical observations. It is reasonable to assume that the decision-maker can access historical observations. In this case, some moment information, such as the mean and variance, may be more readily available than credible parameter value ranges, especially in data-driven paradigms.
As the seminal paper by Scarf et al. (1957) marks the inception, moment-based ambiguity sets have been extensively studied in the literature of RO and DRO for decades (Delage & Ye 2010, Wiesemann et al. 2014, Bertsimas et al. 2019, Ninh 2021), and this concept has also been expanded into into RMDPs. However, it should be noted that the literature on RMDPs employing moment-based ambiguity sets is relatively limited compared to RO and DRO. This scarcity is attributed to the nature of RMDPs, which typically model the environment as dynamics to trigger transitions, without a distinct random variable to provide clear moment information. For instance, the transition probability matrix is defined on a probability simplex in standard RMDPs, complicating the interpretation of mean and variance.
Nevertheless, if the randomness of the decision problem is exogenous, moment-based ambiguity sets can sometimes be tailored well into DRMDPs. For example, in inventory management problems, the demand is often assumed to be an independent random variable concerning initial inventory levels and decisions, and the mean and variance of demand reflect the expected value and fluctuation of uncertainty. However, in the context of DRMDPs, it seems strange to consider transition probabilities as random variables, as their meanings are challenging to decipher in a practical sense. Hence, when detached from specific issues, we suggest carefully considering whether to include moment information into the ambiguity sets of RMDPs with transition uncertainty.
Yang (2018) study the RMDPs where an estimate of the mean and covariance matrix of the disturbance $\bm{\xi}_{t}\sim\mu_{t}$ is the only available information for each stage $t$. Let $\bm{m}_{t}\in\mathbb{R}^{q}$ and $\bm{\Sigma}_{t}\in\mathbb{R}^{q\times q}$ be the estimate of the mean and covariance matrix, respectively. With the stage-wise independent assumption implicitly, the moment-based ambiguity sets are modeled over time as
$\mathcal{P}_{t}\triangleq\left\{\mu_{t}\in\mathfrak{M}(\Xi_{t})~{}\left|~{}\begin{array}[]{l}\mu_{t}(\Xi_{t})=1,\\ |\mathbb{E}_{\mu_{t}}[\bm{\xi}_{t}]-\bm{m}_{t}|\leq\bm{\theta}_{t},\\ \mathbb{E}_{\mu_{t}}[(\bm{\xi}_{t}-\bm{m}_{t})(\bm{\xi}_{t}-\bm{m}_{t})^{\top}]\preceq\gamma_{t}\bm{\Sigma}_{t}\end{array}\right\}\right.,$ (11)
where $\bm{\theta}_{t}\in\mathbb{R}^{k}$ and $\gamma_{t}\geq 1$ are given constants that depend on the confidence in the estimates $\bm{m}_{t}$ and $\bm{\Sigma}_{t}$. From a geometric perspective, the three conditions in (11) denote the following: (i) the support of $\bm{\xi}_{t}$ is $\Xi_{t}$; (ii) the mean of $\bm{\xi}_{t}$ lies in a ball of size $\bm{\theta}_{t}$; and (iii) the centered second-moment matrix of $\bm{\xi}_{t}$ lies in a positive semi-definite cone. Notice that RMDPs with ambiguity set (11) involve an infinite-dimensional mini-max optimization problem. The authors propose a dual reformulation method and prove that there is no duality gap, in virtue of the results in Lasserre (2009). Consequently, the original problem can be transformed into a semi-infinite maximal optimization program. Suppose the objective function is affine, the semi-infinite program in the dual Bellman equation can be shown to be concave with other mild assumptions.
Song et al. (2024) consider a decision-dependent epidemic control problem. Since the transition probability depends on the stochastic epidemiological processes and control manners implemented by the policy maker, the ambiguity sets are endogenous (or decision-dependent), i.e., the transition probability depends on the action $a$ and state $s$. With $(s,a)$-rectangular assumption suggestively, the authors construct ambiguity sets for each $(s,a)$ pair:
$\mathcal{P}_{sa}\triangleq\left\{\mu_{sa}\in\mathfrak{M}_{sa}(\mathcal{S})\left|\mathbf{p}_{sa}\sim\mu_{sa},~{}\bm{\theta}_{sa}^{L}\leq\mathbb{E}_{\mu_{sa}}[\mathbf{p}_{sa}]\leq\bm{\theta}_{sa}^{U}\right\},\right.$ (12)
where $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$ are the lower bound and upper bound of the mean of transition probability. With the ambiguity sets (12), the authors provide the reformulation of Bellman equations. Then, they adopt the linear decision rule to approximate $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$, leading to a mixed integer programming formulation to represent each Bellman equation.
Besides purely utilizing the moment information, Yu & Xu (2015) and Chen et al. (2019) both investigate the generalized-moment-based ambiguity sets, which are also called lifted ambiguity sets (Wiesemann et al. 2014). Generalized-moment-based ambiguity sets involve probability constraints on the support and expectation constraints on the generalized moment of the ambiguous distributions. Before leveraging the powerful modeling capabilities of lifted ambiguity sets, two critical conditions must be met. Firstly, the uncertain parameters across different states must be independent, namely, $s$-rectangular. Secondly, specifically for lifted ambiguity sets, the admissible state-wise ambiguity set $\mathcal{P}_{s}$ for each $s\in\mathcal{S}$ must be represented as the union of marginal distributions across all joint distributions of $(\mathbf{p}_{s},\tilde{n}_{s})$. Here $\tilde{n}_{s}\in\mathcal{N}_{s}$ is a one-dimensional auxiliary random variable that denotes a scenario. Specifically, the RMDPs with lifted ambiguity sets for each state $s\in\mathcal{S}$ can be expressed as follows:
$\mathcal{P}_{s}\triangleq\left\{\mu\in\mathfrak{M}(\mathcal{S}\times\mathcal{N}_{s})\left|\begin{array}[]{lll}&\left(\mathbf{p}_{s},\tilde{n}_{s}\right)\sim\mu\\ \vspace{1ex}&\mathbb{E}_{\mu}\left.\left[\mathbf{p}_{s}~{}\right|~{}\tilde{n}_{s}\in\mathcal{N}_{j}\right]=\theta_{j}&\forall j\in\left[J_{s}\right]\\ \vspace{1ex}&\mu\left[\mathbf{p}_{s}\in\mathcal{D}_{n}~{}|~{}\tilde{n}_{s}=n\right]=1&\forall n\in\mathcal{N}_{s}\\ &\mu\left[\tilde{n}_{s}=n\right]=\omega_{n}&\forall n\in\mathcal{N}_{s}\\ &\text{for some }\bm{\omega}\in\Delta^{|\mathcal{N}_{s}|},\left(\theta_{j},\beta_{j}\right)\in\mathcal{U}_{j}&\forall j\in\left[J_{s}\right]\end{array}\right\}\right.\end{aligned},$ (13)
where for each $j\in[J_{s}]$, the set $\mathcal{N}_{j}\subseteq\mathcal{N}_{s}$ is a subset of scenarios, and sets $\mathcal{D}_{n}\subseteq\Delta^{|\mathcal{S}|}$.
@@
$\mathbb{E}_{\mu}\left.\left[g_{j\tilde{n}_{s}}(\mathbf{p}_{s})~{}\right|~{}\tilde{n}_{s}\in\mathcal{N}_{j}\right]\leq\beta_{j}\quad\forall j\in\left[J_{s}\right],$
@@
where the function $g_{j\tilde{n}_{s}}$ is convex lower semi-continuous, and the authors explore and formulate $g_{j\tilde{n}_{s}}$ as the Wasserstein distance. Additionally, they show that this enhanced formulation can be effectively modeled by algebraic modeling packages (Chen et al. 2020) and solved by off-the-shelf commercial solvers.

### v4

@@ 3 Parametric RMDPs
where $\Theta$ is a subset of $\mathbb{R}^{q}$ and $P^{\theta}(\cdot\mid s,a),~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ is an affine function from $\Theta$ to $[0,1]$ that satisfies $P^{\theta}(\cdot\mid s,a):=k_{sa}+K_{sa}\theta$ for some $k_{sa}\in\mathbb{R}^{|\mathcal{S}|}$ and $K_{sa}\in\mathbb{R}^{|\mathcal{S}|\times k}$. The affine function form implies that the conditional transition probability is the outcome of a linear regression, where the coefficients are $(K_{sa},k_{sa})$ regarding the parameters $\theta$. Note that $\theta$ controls the ambiguity sets essentially, which allows to condense all ambiguous information and parameters in the set $\Theta$, e.g.,
@@
where $O_{l}$ is a $k\times k$ matrix, $O_{l}\preceq 0$, $\omega$ is the constant and $L$ is the number of constraints applied to $\theta$. We assume that $\Theta$ has a nonempty interior, that is, none of the parameters in $\Theta$ is fully explained by the others. The benefit of this construction defined in (14) is that the quadratic inequalities in $\Theta$ with negative semidefinite matrices can be reformulated as second-order cone constraints, allowing RMDPs to be expressed as second-order cone programs (SOCPs) or semidefinite programs (SDPs), which are convex problems and can be solved efficiently with interior-point methods. The authors assume that $\Theta$ is bounded and has a nonempty interior, which means none of the parameters in $\Theta$ is fully explained by the others. With the $s$-rectangular assumption, they show that the robust Bellman Optimality equation holds with the optimization procedure proceeds for each state $s\in\mathcal{S}$ and such operators are contraction mappings, ensuring the unique optimal value-to-go function. Furthermore, the authors provide the time complexity of policy evaluation and policy improvement routines, respectively.
@@
Black et al. (2023) apply RMDPs for newsvendor problems, where the demand is exogenous and state transitions are essentially driven by demand realizations (underlying dynamics). Thus, the transition probability matrix is not time-invariant, and given the inventory position, the conditional transition probabilities can be equivalently derived by the distribution of demands. In this case, the authors investigate RMDPs where the ambiguity sets are limited to distributions within the same parametric family, focusing on identifying the worst-case parameters rather than the entire distribution. This approach aims to alleviate challenges with moment estimation in non-parametric RMDPs, as well as the issue of overly conservative solutions of non-parametric methods that can arise when parametric distributions are fitted well enough. Let $f_{\xi_{sa}}$ be the probability mass function of exogenous random variable $\xi_{sa}$ which is controlled by the parameters $\bm{\theta}_{sa}=(\theta_{sa_{1}},\cdots,\theta_{sa_{k}})\in\mathbb{R}^{k}$. The authors assume that the next state $s_{t+1}$ is specified by a simple, known function $g$ of $\xi_{sa}$ as $s_{t+1}=g(\xi_{sa}|s,a)$. For example, in inventory management, states often denote the initial inventory positions and actions are the order quantities. Consequently, one possible function is $g=s+a-\hat{\xi}_{sa}$ with the demand realization $\hat{\xi}_{sa}$. As we discussed in REMARK 2.1, for each $(s,a)\in\mathcal{S}\times\mathcal{A}$, conditional transition (discrete) distribution $\mathbf{p}_{sa}$ can be uniquely represented as the distribution of $\xi_{sa}\in\Xi_{sa}$. More precisely, each transition probability $p_{sas^{\prime}}$ for the transition tuple $(s,a,s^{\prime})$ can be computed as
$p_{sas^{\prime}}=P^{\bm{\theta}_{sa}}(s^{\prime}\,|\,s,a)=P^{\bm{\theta}_{sa}}(g(\xi_{sa}\,|\,s,a)=s^{\prime})=\sum_{\xi\in\Xi_{sa}(s^{\prime})}f_{\xi_{sa}}(\xi\,|\,\bm{\theta}_{sa}),$ (15)
@@
where $\Xi_{sa}(s^{\prime})$ is the support of $\xi$ with $(s,a,s^{\prime})$ transition. Because the transition probability matrix $\mathbf{p}$ is uniquely specified by $\bm{\theta}$, the authors simplify ambiguity sets for $\bm{\theta}$, rather than $\mathbf{p}$, to maintain the equivalence structure as in (15). They consider ambiguity sets with $s$-rectangular structure $\Theta=\bigotimes_{s\in\mathcal{S}}\Theta_{s}$ where $\Theta_{s}\subseteq\mathbb{R}^{k},\mathrm{~{}}\forall s\in\mathcal{S}$, and then reformulate the RMDP as:
$\max_{\pi}\min_{\bm{\theta}\in\Theta}\mathbb{E}_{\mathbf{p}}^{\pi}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\,\right|\,\bm{\theta}\right].$ (16)
In this paper, the authors also demonstrate how to construct the uncertainty set in terms of $\bm{\theta}$ when finite samples from the true distribution of $\xi$ are accessible. By standard results in maximum likelihood theory (Millar 2011), given the log-likelihood function $\ell(\cdot)$ for observed data, a Maximum Likelihood Estimation (MLE) $\hat{\bm{\theta}}_{sa}$ of true parameter $\bm{\theta}^{0}_{sa}$ satisfies:
$\sum_{a\in\mathcal{A}}\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}^{0}\right)^{T}I_{\mathbb{E}}\left(\bm{\theta}_{sa}^{0}\right)\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}^{0}\right)\thicksim\chi_{k|\mathcal{A}|}^{2},$
where $I_{\mathbb{E}}\left(\bm{\theta}^{0}_{sa}\right)=\left(-\mathbb{E}_{\xi_{sa}}\left[\frac{\partial}{\partial\theta^{0}_{sai}\partial\theta^{0}_{saj}}\ell({\bm{\theta}}^{0}_{sa})\right]\right)_{i,j=1,...,k}$ is the expected Fisher information matrix and $\chi_{k|\mathcal{A}|}^{2}$ is the chi-squared distribution with $k|\mathcal{A}|$ degrees of freedom. As the asymptotic equivalence holds, the authors use $\hat{\bm{\theta}}_{sa}$ estimated via historical data to approximate $\bm{\theta}^{0}_{sa}$ and construct a $1-\alpha$ confidence level set given by:
@@
$\Theta_{s}=\left\{\bm{\theta}_{s}\in\mathbb{R}^{|\mathcal{A}|}\times\mathbb{R}^{k}\,\middle|\,\sum_{a\in\mathcal{A}}\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)^{T}I_{\mathbb{E}}\left(\hat{\bm{\theta}}_{sa}\right)\left(\hat{\bm{\theta}}_{sa}-\bm{\theta}_{sa}\right)\leq\chi_{k|\mathcal{A}|,1-\alpha}^{2}\right\}\quad\forall s\in\mathcal{S}.$ (17)
The uncertainty set $\Theta_{s}$ contains all parameter vectors $\bm{\theta}_{s}$ that are “close enough” to the MLE estimate, where “close enough” is defined such that the sum of the distances (i.e., the quadratic term weighted by the Fisher Information matrix) over all actions must be less than or equal to the $(1-\alpha)$ quantile of $\chi_{k|\mathcal{A}|}^{2}$. Considering the non-linearity of $\mathbf{p}$ as functions of the parameters, they replace the uncertainty set $\Theta_{s}$ with a discretized set $\Theta_{s}^{\prime}$ and reformulate the Bellman update (3) as a linear program with $|\Theta_{s}^{\prime}|+1$ constraints.
@@ 4 Moment-Based RMDPs
The motivations behind parametric RMDPs are relatively intuitive, however, determining reasonable ranges for parameter values in practice remains a significant challenge. Moreover, most parametric approaches intend to impose constraints on each transition probability $p_{sas^{\prime}}$ such that the uncertainty sets are excessively large. While this ensures general applicability, it often leads to overly conservative or impractical solutions. On the other hand, it is more practical to assume that decision-makers have confidence in some probabilistic information regarding the dynamics. Under such circumstances, moments of the distribution, such as the mean and variance, are typically more accessible and reliable than precise parameter ranges, as they can be readily inferred from historical data.
@@
Nevertheless, moment-based ambiguity sets can sometimes be tailored well to RMDPs where the randomness of the decision problem is exogenous, namely, unrelated to states and actions. It is an interesting concept discussed as exogenous process recently in model-based RL algorithms (Madeka et al. 2022, Sinclair et al. 2023). Besides inventory management discussed above, numerous applications exhibit characteristics of exogenous processes, including appointment scheduling, airline revenue management, and resource allocation. While it may not be natural to understand and estimate the moments of transition probabilities in general RMDPs or DRMDPs, it is important to carefully include the moment-based information in the construction of ambiguity sets in many applications.
Yang (2018) study the RMDPs where the transition is controlled by state $s$, action $a$, stochastic disturbance $\bm{\xi}$, and a measurable function $f$  [^1] More precisely, this formulation is more related to the concept of optimal control (OC). However, under mild assumptions and appropriate settings, we can reformulate an OC as an MDP, regarding significant overlap in the key techniques utilized by both methodologies.:
$s_{t+1}=f(s_{t},a_{t},\bm{\xi}_{t})\quad\forall t\in[T].$
In this case, the author supposes that estimates of the mean, $\bm{m}_{t}\in\mathbb{R}^{q}$, and covariance matrix, $\bm{\Sigma}_{t}\in\mathbb{R}^{q\times q}$, of the disturbance $\bm{\xi}_{t}\sim\mu_{t}$ are the only available information for each stage $t$. With the stage-wise independent assumption implicitly, the moment-based ambiguity sets are modeled over time as
$\mathcal{P}_{t}\triangleq\left\{\mu_{t}\in\mathfrak{M}(\Xi_{t},\mathcal{F}^{t})~{}\left|~{}\begin{array}[]{l}\mu_{t}(\Xi_{t})=1\\ |\mathbb{E}_{\mu_{t}}[\bm{\xi}_{t}]-\bm{m}_{t}|\leq\bm{\theta}_{t}\\ \mathbb{E}_{\mu_{t}}[(\bm{\xi}_{t}-\bm{m}_{t})(\bm{\xi}_{t}-\bm{m}_{t})^{\top}]\preceq\beta_{t}\bm{\Sigma}_{t}\end{array}\right\}\right.,$ (18)
where $\bm{\theta}_{t}\in\mathbb{R}^{k}$ and $\beta_{t}\geq 1$ are given constants that depend on the confidence in the estimates $\bm{m}_{t}$ and $\bm{\Sigma}_{t}$. From a geometric perspective, the three constraints in (18) denote the following: (i) the support of $\bm{\xi}_{t}$ is $\Xi_{t}$; (ii) the mean of $\bm{\xi}_{t}$ lies in a ball of size $\bm{\theta}_{t}$; and (iii) the centered second-moment matrix of $\bm{\xi}_{t}$ lies in a positive semidefinite cone. Thus, this ambiguity models how likely $\bm{\xi}_{t}$ is to be close to the estimate $\bm{m}_{t}$ in terms of the weighted correlation matrix estimate ${\beta_{t}}\bm{\Sigma}_{t}$. The parameters $\bm{\theta}_{t}$ and ${\beta_{t}}$ allow for adjusting the size of the ambiguity set based on the confidence in the estimates.
Notice that solving an RMDP with an ambiguity set (18) involves an infinite-dimensional maximin optimization problem, which is generally computationally intractable. To overcome the challenge, the author proposes a dual reformulation of the inner minimization problem, which is a semi-infinite maximal optimization program. Leveraging the results from Lasserre (2009), it is shown that no duality gap exists in this reformulation. Consequently, by substituting the inner problem in the Bellman equation without sacrificing optimality, the author develops the dual Bellman equation which can be shown to be concave with some mild assumptions, e.g., measurable function $f$ is affine, and state space $\Xi_{t}$ is convex and compact.
Song et al. (2024) consider a decision-dependent epidemic control problem where the transition probabilities depend not only on the stochastic epidemiological processes but also on control manners implemented by the policy-maker. Thus, the ambiguity sets are endogenous (or decision-dependent), i.e., the transition probabilities depend on the action $a$ and state $s$. With $(s,a)$-rectangular assumption implicitly, the authors construct ambiguity sets for each $(s,a)$ pair:
@@
$\mathcal{P}_{sa}\triangleq\left\{\mu_{sa}\in\mathfrak{M}(\mathcal{S},\mathcal{F}^{\mathcal{S}})\mid\mathbf{p}_{sa}\sim\mu_{sa},~{}\bm{\theta}_{sa}^{L}\leq\mathbb{E}_{\mu_{sa}}[\mathbf{p}_{sa}]\leq\bm{\theta}_{sa}^{U}\right\},$ (19)
where $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$ are the lower bound and upper bound vectors of the mean of transition probabilities, and $\mathcal{F}^{\mathcal{S}}$ is the corresponding $\sigma$-algebra when sample space is $\mathcal{S}$. Facing the same issue that RMDPs with moment-based ambiguity set are generally intractable, Song et al. (2024) first relax the hard constraint $\bm{\theta}_{sa}^{L}\leq\mathbb{E}_{\mu_{sa}}[\mathbf{p}_{sa}]\leq\bm{\theta}_{sa}^{U}$ into soft constraints:
$\int\mathbf{p}_{sa}\mathrm{d}\mu_{sa}(\mathbf{p}_{sa})-\bm{\theta}_{sa}^{U}\leq\bm{x},\quad\bm{\theta}_{sa}^{L}-\int\mathbf{p}_{sa}\mathrm{d}\mu_{sa}(\mathbf{p}_{sa})\leq\bm{x},$
and adjust the objective function (i.e., the Bellman equation) by penalizing constraint violations with $k\mathbf{1}^{T}\bm{x}$, where $k$ represents a user-specified penalty coefficient. Subsequently, the authors apply the standard Lagrangian dualization approach to obtain the dual Bellman equation. Slightly different from direct dualization, penalty coefficient $k$ offers greater flexibility and allows the incorporation of expert knowledge into the model. Furthermore, the authors consider $\bm{\theta}_{sa}^{L}$ and $\bm{\theta}_{sa}^{U}$ can be expressed by the linear functions of action $a$, e.g.,
$\theta_{sa}^{U}(s^{\prime})=\rho_{0}(s^{\prime})+\sum_{i=1}^{N_{a}}\rho_{i}^{s}(s^{\prime})a_{i}\quad\forall s^{\prime}\in\mathcal{S},$
$\theta_{sa}^{L}(s^{\prime})=\eta_{0}(s^{\prime})+\sum_{i=1}^{N_{a}}\eta_{i}^{s}(s^{\prime})a_{i}\quad\forall s^{\prime}\in\mathcal{S},$
where $a_{i}$ denotes the $i$-th dimension of action $a$ and $N_{a}$ is the total dimension. Coefficients $\bm{\rho}^{s}$ and $\bm{\eta}^{s}$ can be obtained by linear regression or machine learning approaches, and $\rho_{i}^{s}(s^{\prime})$ (or $\eta_{i}^{s}(s^{\prime})$) is the coefficient of transition $(s,a,s^{\prime})$ where $i$-th dimension of action $a$ is $a_{i}$. Although the linear decision rule overcomes the challenge of dimensionality, it introduces the bilinear terms in the dual Bellman equation (i.e., $\bm{\theta}_{sa}^{U}$ or $\bm{\theta}_{sa}^{L}$ with the Lagrangian multipliers). The authors adopt McCormick envelope relaxation (McCormick 1976) and exact unary expansion (Gupte et al. 2013), leading to a mixed integer programming formulation to represent each Bellman equation.
Rather than purely utilizing the moment information, Yu and Xu (2015) and Chen et al. (2019) both investigate the generalized-moment-based ambiguity sets, which are also called lifted ambiguity sets (Wiesemann et al. 2014). Generalized-moment-based ambiguity sets involve probability constraints on the support and expectation constraints on the generalized moment of the ambiguous distributions. Before leveraging the powerful modeling capabilities of lifted ambiguity sets, two critical conditions must be met. Firstly, the uncertain parameters across different states must be independent, namely, $s$-rectangular. Secondly, the admissible state-wise ambiguity set $\mathcal{P}_{s}$ for each $s\in\mathcal{S}$ must be representable as the union of marginal distributions across all joint distributions of $(\mathbf{p}_{s},\tilde{n}_{s})$. Here $\tilde{n}_{s}\in\mathcal{N}_{s}$ is a one-dimensional auxiliary random variable that denotes a scenario. In essence, we could view $\tilde{\mathcal{S}}=\mathcal{S}\times\mathcal{N}_{s}$ as the new state space in which the original state space incorporates the scenario information.
In Yu and Xu (2015), the RMDPs with lifted ambiguity sets for each state $s\in\mathcal{S}$ can be expressed as follows:
$\mathcal{P}_{s}\triangleq\left\{\mu\in\mathfrak{M}(\tilde{\mathcal{S}},\mathcal{F}^{s})\left|\begin{array}[]{lll}&\left(\mathbf{p}_{s},\tilde{n}_{s}\right)\sim\mu\\ \vspace{1ex}&\mathbb{E}_{\mu}\left[\mathbf{p}_{s}\mid\tilde{n}_{s}\in\mathcal{N}_{j}\right]=\theta_{j}&\forall j\in\left[J_{s}\right]\\ \vspace{1ex}&\mu\left[\mathbf{p}_{s}\in\mathcal{D}_{n}\mid\tilde{n}_{s}=n\right]=1&\forall n\in\mathcal{N}_{s}\\ &\mu\left[\tilde{n}_{s}=n\right]=\omega_{n}\ \text{for some }\bm{\omega}\in\Delta^{|\mathcal{N}_{s}|}&\forall n\in\mathcal{N}_{s}\\ &\theta_{j}\in\mathcal{U}_{j}&\forall j\in\left[J_{s}\right]\end{array}\right\}\right.\end{aligned},$ (20)
where for each $j\in[J_{s}]$, the set $\mathcal{N}_{j}\subseteq\mathcal{N}_{s}$ is a subset of scenarios, and sets $\mathcal{D}_{n}\subseteq\Delta^{|\mathcal{S}|}$. The notion of $J_{s}$ is the cardinality of possible scenarios with state $s$. The constraints in (20) denote the following: ($i$) conditional mean of $\mathbf{p}_{s}$ is $\theta_{j}$ in terms of $\tilde{n}_{s}\in\mathcal{N}_{j}$; ($ii$) support constraints $\mathbf{p}_{s}\in\mathcal{D}_{n}$ for each scenario $\tilde{n}_{s}=n$; ($iii$) probability mass (or weight) for each scenarios, where $\bm{\omega}$ is a probability simplex; and ($iv$) additional constraints on parameters.
[…]
$\mathbb{E}_{\mu}\left.\left[g_{j\tilde{n}_{s}}(\mathbf{p}_{s})~{}\right|~{}\tilde{n}_{s}\in\mathcal{N}_{j}\right]\leq\beta_{j}\quad\forall j\in\left[J_{s}\right],$ (21)
and replacing parameter constraints $\{\theta_{j}\in\mathcal{U}_{j}\ \forall j\in\left[J_{s}\right]\}$ with
$(\theta_{j},\beta_{j})\in\mathcal{U}_{j}\quad\forall j\in\left[J_{s}\right].$ (22)
@@
The function $g_{j\tilde{n}_{s}}$ in (21) is convex lower semi-continuous, and the authors explore and formulate $g_{j\tilde{n}_{s}}$ as the Wasserstein distance. Additionally, they show that this enhanced formulation can be effectively modeled by algebraic modeling packages (Chen et al., 2020) and solved by off-the-shelf commercial solvers.